# The Yōkai Learning Environment: Tracking Beliefs Over Space and Time

### v1

@@ 6 Related Work
Current benchmarks study ToM reasoning in AI mostly from a static observer’s perspective Baker et al. (2011, 2017); Rabinowitz et al. (2018); Bara et al. (2021, 2023); Fan et al. (2021); Duan et al. (2022); Kim et al. (2023); Bortoletto et al. (2024c); Gandhi et al. (2023), framed as a supervised learning task – an approach with notable limitations Aru et al. (2023); Bortoletto et al. (2024a). Other benchmarks evaluated agents in interactive environments, such as the Hanabi Learning Environment (HLE) Bard et al. (2020) and SymmToM Sclar et al. (2022). In contrast to Hanabi, Yōkai requires spatial-temporal belief reasoning, requiring agents to bind beliefs to moving cards Fernandez et al. (2023). Compared to SymmToM, YLE offers greater strategic depth and stronger pressure to develop ToM. In SymmToM, heuristics outperform all baselines and communication is unrestricted, while in YLE, misused hints can irreversibly lock cards in losing positions. Unlike Hanabi and SymmToM, YLE also introduces a high-stakes belief-based decision: agents may end the game early for a higher payoff, sacrificing future information and relying more on their beliefs.
@@
A second line of work investigated whether neural networks represent mental states internally Bortoletto et al. (2024a); Zhu et al. (2024). For example, Matiisen et al. Matiisen et al. (2022) probed DRL agents for intention representations in cooperative navigation. We extend probing to YLE, where beliefs evolve dynamically and asymmetric knowledge emerges over time – showing the value of probing in a more complex environment.
Finally, YLE contributes to the broader family of cooperative multi-agent RL (MARL) benchmarks Samvelyan et al. (2019); Carroll et al. (2019); Bard et al. (2020); Ruhdorfer et al. (2024), but with a strong emphasis on ToM and common ground reasoning. While many techniques were developed for Hanabi, our results indicate they do not transfer directly to the YLE, thus highlighting the new challenges posed by our benchmark and requiring new modelling approaches.
7 Conclusion & Future Work
We introduced the Yōkai Learning Environment – a multi-agent RL benchmark where effective collaboration requires belief tracking over space and time. Implemented in JAX, YLE supports end-to-end GPU training at hundreds of thousands of steps per second. We show that RL agents on the YLE struggle to retain knowledge, to generalise across partners, or to identify when coordination has succeeded – a high-stakes decision reflected in their inability to end games early. YLE thus offers a scalable, diagnostic testbed for advancing common ground reasoning in collaborative AI.
While our work focused on evaluating RL agents, future work might use this benchmark to evaluate other systems, like language agents Xi et al. (2023); Wang et al. (2024). We envision that Yōkai can play a major role in isolating and assessing critical aspects of these systems that realte to memory, spatial reasoning and common ground tracking.
Acknowledgements
@@
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting C. Ruhdorfer. We especially thank R. Yang for implementing a prototype of the environment as part of her master’s thesis, and A. Penzkofer and L. Shi for numerous insightful discussions.
Appendix A Appendix: Infrastructure & Tools
Servers and software
@@ Appendix A Appendix: Infrastructure & Tools
We ran our experiments on two different server systems. Both are running Ubuntu 24.04. One is equipped with NVIDIA Tesla V100-SXM2 GPUs with 32GB of memory and Intel Xeon Platinum 8260 CPUs. The other is equipped with NVIDIA H100-NVL GPUs with 94GB of memory and AMD EPYC 9454 CPUs. All training runs are executed on a single GPU only. We trained our models using JAX Bradbury et al. [2018], Flax Heek et al. [2023] and Optax DeepMind et al. [2020]. In terms of tools, we did our analysis using NumPy Harris et al. [2020], Pandas Team [2020], SciPy Virtanen et al. [2020] and Matplotlib Hunter [2007]. Our TransformerXL implementation is based on the one provided by Hamon [2024]. Our single-file implementation of MAPPO is based on the one provided by JaxMARL Rutherford et al. [2023] with additional inspiration taken from Hamon [2024].
Accesability and reusability
@@
The Yōkai Learning Environment (YLE) integrates directly into JaxMARL and implements its standard environment interface. As a result, any baseline method implementation of JaxMARL can be adapted to our environment quickly and with minimal effort. This makes our open-source environment easily accessible to future researchers who will benefit from our reference implementations and those provided by the wider community. It also enables inter-environment comparison, facilitating broader benchmarking across cooperative multi-agent tasks. The project is accessible under https://git.hcics.simtech.uni-stuttgart.de/public-projects/Yokai_env.
Training times and compute requirements
Training times vary between server setup and agent. Our longest training agent is ToMTrack-OP+C+R since it needs to compute memory updates for the TransformerXL and the asymmetric domain randomisation used in Other-Play. Experiments with ToMTrack-OP+C+R take 8 hours to complete per seed on our H100 setup and around 24 hours on our V100 setup. The CNN-M baselines train quickest:  6 hours on the H100 server and  14 hours on a V100 server. All our experiments are runnable on a single V100 with 32GB VRAM.
Appendix B Fast, Scalable, and Parallel Yōkai in Graph-based JAX
@@ B.1 JAX-based Implementation
Resulting Performance
Num Envs $n$ 512 1,024 2,048
2-player 3x3 139,107 274,669 559,251
2-player 4x4 135,662 144,636 148,269
Table 3: We evaluate steps-per-second (SPS) of our environment on a single Nvidia H100 by taking random actions in $n$ parallel environments. Our environment scales to hundreds of thousands of SPS.
We present how our environment performs in terms of steps-per-second across many parallel environments. All experiments were run by taking 1000 random consecutive actions in $n$ parallel environments. Table 3 shows results. Our environment scales similarly to other recently proposed JAX-based RL environments, e.g. Rutherford et al. [2023]; Nikulin et al. [2023]; Ruhdorfer et al. [2024]; Matthews et al. [2024].
B.2 Reinforcement Learning in the YLE
@@
While in the original game, cards can be placed in any legal configuration, our environment limits the configuration to be in a predefined field of size $g\times g$ where $g=9$ in the 3x3 and $g=10$ in the 4x4 version. This enables the efficient computation of legal moves but prohibits some rare card configurations. While computing the exact number of disallowed configurations is infeasible, we find them to be extremely uncommon in practice.
For example, let’s assume a 2-player 3x3 game on a 9x9 grid. In such a setting, there a 4 hints and thus a total of 8 move actions for both players together. Cards are initially placed in a square in the centre of the board, leaving 3 empty spaces in the grid towards either side. In theory, thus, if agents agree to start placing cards in a row towards one of these sides, they would run out of space towards the border after 3 moves. In practice, this is seldom a suitable strategy and thus does not usually occur in games of Yōkai. This is because players want to group cards and are limited by the number of moves. They thus do not tend to start grouping cards in random directions, but rather group cards where they initially find a certain colour.
To verify that this is also true for our agents, we count the number of times when, in an evaluation game, an agent places a card on a border position. Note this is an overestimation as it does not account for the cases where agents place a card on the border but do not intend to place further cards in the same direction. We find that policies rarely to never use to border region of the board. Our best-performing policy ToMTrack uses the border $0.0\pm 0.0\%$ of the time when evaluated over all three seeds in $1000$ games each. We also discuss this in section 5.7.
@@
For example, the action space is of size $1,068$ for 2-player 3x3 games and $2,322$ for the 4x4 version.
@@
Sclar et al. [2022] defines symmetric ToM environments as environments that (1) have a symmetric action space, (2) feature imperfect information, (3) allow the observation of others, and (4) require information-seeking behaviour. YLE features all four and thus qualifies. First, the action space is symmetric (see above). Second, information is imperfect since agents can never observe all cards first-hand and, for most of the game, only know a very limited subset of all cards. Third, it allows the observation of others both using explicitly encoded actions, as in ToMTrack or by observing their moves indirectly via their observations. Fourth, information needs to be acquired by observing cards first-hand or, second-hand, by interpreting moves and reasoning about knowledge. Agents that do not seek information and, for instance, repentantly observe the same card, will fail the game.
@@
4x4 3x3
@@
Table 4: Number of hint cards according to the number of players and the environment version. The 3x3 version features three colours and thus has no three-colour hints.
How many and which hint cards are available depend on the environment version and the number of players. For a 2-player game, there are 2 hints of one, 3 hints of two and 2 hints of three colours, i.e. 7 in total. Since there are 4 possible one-, 6 possible two- and 4 possible three-colour hints, players get a random subset of these hints at the start of the game per category. We give a full overview of all possible variations in Table 4. The number of hints directly influences the maximum number of turns in the game. This is because the game ends when all hints have been used. This means that games take longer if more players are added. Note, though, that this usually does not make the game easier. Instead, the number of cards one player can directly observe themself decreases with the number of players. For 4x4 YLE in the 2-player version, any agent can observe up to 14 of the 16 cars themselves (8 of 9 in 3x3 Yōkai). This shrinks to 10 of 16 for the 4x4 version and 6 of 9 for the 3x3 version when 4 players take part. Note how in the 3- and 4-player versions, agents are thus required to both work with less first-hand information and need to rely more on hints and interpreting the moves made by those around them. Not only this, but agents are tasked to do this tracking for more players as well. Thus, adding players to the YLE increases the difficulty substantially in two ways.
@@
B.3 Second Order ToM Reasoning Example
Figure 12: A second-order ToM reasoning example over multiple timesteps in 2-player YLE. Refer to caption beliefs that Refer to caption beliefs that he (Refer to caption ) knew that cards 1 and 2 are of the same colour. Even though Refer to caption never saw card 1 and Refer to caption never observed card 3, both now know where all blue cards are, that they are grouped and that both share this knowledge as part of their common ground. In the future, they can now both observe and move different cards to finish early as they have one less card that needs to be observed.
Figure 12 displays an example for second-order ToM reasoning in 2-player Yokai. It shows an agent ([Uncaptioned image] ) making inferences about what they assume the other agents ([Uncaptioned image] ) knowledge is due to their own previous action. In the example [Uncaptioned image] infers the colour of the card due to the behaviour of the other agent. [Uncaptioned image] reasons that because of how [Uncaptioned image] played, [Uncaptioned image] must have interpreted his move correctly. [Uncaptioned image] thus reasons that [Uncaptioned image] thinks that [Uncaptioned image] knew that cards 1 and 2 have the same colour and must go together. This is the case, [Uncaptioned image] made this inference correctly, and card 3 indeed is also blue. This reasoning allows agents to require less first-hand observations to get an overview of the board, potentially enabling them to end early.
@@ B.3 Second Order ToM Reasoning Example
Appendix C Training Details
We share all the details of the training here, including all hyperparameters used in training and all details on the network architectures and how we tuned them. In general, we have tuned many aspects of our policy to arrive at the strongest agents we could for our experiments, see below. Finally, the number of parameters per model is shown in Table 6. We took extra care in making these equal to make our results comparable.
C.1 Reinforcement Learning Details
@@
Optimiser Adam Kingma and Ba [2015]
@@
Table 5: Hyperparameters of the learning process.
Our work employs a standard MAPPO algorithm Yu et al. [2022] for all experiments. MAPPO is an extension of PPO Schulman et al. [2017] to the multi-agent setting. It separates the actor and the critic network. Since the actor and critic are separated and the critic is discarded after training, MAPPO allows giving the critic access to privileged information during training. This leads to the good performance of MAPPO in cooperative multi-agent tasks, compare Yu et al. [2022], which is why we choose it in our work. We present all parameters of the learning process in Table 5.
C.2 Neural Network Architectures
Model # Parameters
GCN-M 317,048
GATv2-M 317,560
RM-M 354,168
CNN-M 322,156
CNN-LSTM 470,252
CNN-GRU 437,484
CNN-S5 388,524
CNN-TrXL 400,940
ToMTrack 404423
ToMTrack-SAD 417863
Table 6: Number of parameters of all models used in this work.
Our work employs several encoder architectures. We present all the details on them below. In general, we aimed to make comparisons fair by giving all compared models a similar number of parameters, also see Table 6. The biggest advantage in terms of parameter size was RM-M for the spatial reasoning and CNN-LSTM for the temporal reasoning experiment, but neither performed the best.
C.2.1 General Policy Architecture
In general, in our work, no matter the encoder, the final policy network always has the same structure. That is given an embedding $e_{\text{encoder}}$ produced by an encoder network $f_{\text{encoder}}$ and possibly a memory architecture, we pass it to a stack of dense layers and either a value or action head. We always use $4$ dense layers before the heads with ReLU Fukushima [1975]; Nair and Hinton [2010] as the activation function between the layers and a hidden dimension of $128$. The layers are initialised using orthogonal initialisation with a scale of $\sqrt{2}$ and bias 0. We have determined the number of policy layers and their size using hyperparameter tuning. We compared $2$, 33 and $4$ layers with hidden dimension $32$, $64$ and $128$. We found that the largest configuration outperforms the others. Additionally, for all neural network architectures detailed below, we also varied and explored: The number of encoder blocks (2, 3 and 4), the hidden dimension of the encoder blocks, the aggregation function after the encoder (mean, sum, multiply and concatenate) and whether to anneal the dense rewards (False and True). Finally, the value head produces a singular scalar output and is initialised using orthogonal initialisation with a scale of $0.00$ instead of $\sqrt{2}$. We initialise the action head in the same way but it produces logits for all possible actions instead. The critic network is always parameterised the same as the actor network, but the input is different. Since we are using MAPPO our critic receives as input a so-called world state that combines the individual observations of all agents. This is possible in MAPPO due to centralised training and decentralised execution Foerster et al. [2016]. As the world state, we concatenate the individual observations along the feature dimension.
C.2.2 GCN, GATv2 and the pre-embedding of Nodes

### v2

@@ 8 Related Work
Current benchmarks study ToM reasoning in AI primarily from a static observer’s perspective (Baker et al., 2011; 2017; Rabinowitz et al., 2018; Bara et al., 2021; 2023; Fan et al., 2021; Duan et al., 2022; Kim et al., 2023; Bortoletto et al., 2024b; Gandhi et al., 2023; Bortoletto et al., 2025), often framed as supervised learning problems (Aru et al., 2023; Bortoletto et al., 2024a). In contrast, interactive benchmarks such as the Hanabi (Bard et al., 2020) and SymmToM (Sclar et al., 2022) evaluate ToM in multi-agent interactive settings. YLE extends interactive ToM benchmarks by introducing dynamic entity movement, ambiguous communication, and early termination decisions within a cooperative Dec-POMDP. Yōkai was first explored in AI research in the context of formal logics (Fernandez et al., 2023), and, additionally, has been explored in a preliminary study on the feasibility of training self-play RL agents for Yōkai (Yang, 2024). The present work instead introduces the YLE as an open-source benchmark for belief tracking and zero-shot coordination.
@@
A second line of work investigates whether neural networks internally represent mental states (Bortoletto et al., 2024a; Zhu et al., 2024). For example, Matiisen et al. (2022) probe RL agents for intention representations in cooperative navigation. We extend probing analyses to YLE, where hidden representations must encode evolving beliefs under partial observability.
Alongside YLE, OvercookedV2 (Gessler et al., 2025) has recently been proposed as a benchmark for ZSC, augmenting the Overcooked environment (Carroll et al., 2019) with partial observability and asymmetric information. YLE differs in that it introduces substantially larger state and action spaces and, in configurations with more than three players or 16 Yōkai cards, poses a significant challenge even in SP. Additionally, YLE offers a full reference implementation of OBL, which previously achieved state-of-the-art inter-seed cross-play performance in the HLE. Our results show that methods with strong performance in Hanabi perform substantially worse in YLE and that algorithm rankings change, highlighting the importance of diverse benchmarks for evaluating ZSC algorithms. YLE also contributes to the broader family of cooperative multi-agent RL benchmarks (Samvelyan et al., 2019; Ruhdorfer et al., 2025b; Tomilin et al., 2025), with a particular emphasis on belief tracking, ToM, and common ground formation.
Finally, we consider the ZSC setting as introduced by Hu et al. (2020) and further clarified in (Treutlein et al., 2021). Other broader notions of coordination with novel partners exist and are often studied under the label of ad-hoc teamwork (AHT) (Stone et al., 2010). Representative AHT methods are typically population- and diversity-based and include (Strouse et al., 2021; Lucas & Allen, 2022; Xue et al., 2022; Zhao et al., 2023; Yan et al., 2023; Hui et al., 2025; Ruhdorfer et al., 2025a). In this work, we focus on ZSC first to isolate the coordination challenges posed by the YLE, reasoning that AHT progress on the YLE is only possible after making progress on the ZSC problem. We thus view evaluating AHT methods in the YLE as valuable future work after ZSC performance has improved, similar to the progression observed in Hanabi research (Dizdarevic et al., 2025).
9 Conclusion
We introduced the Yōkai Learning Environment – a new multi-agent RL benchmark in which effective collaboration requires tracking and maintaining beliefs over space and time. Implemented in JAX, YLE supports end-to-end GPU training at hundreds of thousands of steps per second. Empirically, we show that state-of-the-art ZSC methods that achieve near-perfect inter-seed cross-play in Hanabi exhibit persistent SP-XP gaps in YLE, including degraded early-ending frequency and calibration. These results indicate that coordination in YLE places additional structural demands on belief formation and common ground reasoning and that apparent ZSC progress may be benchmark-specific. The YLE thus provides a complementary benchmark for evaluating ZSC beyond current environments.
Broader Impact Statement
This work introduces a new environment for studying fundamental aspects of cooperation in reinforcement learning agents. As such, it supports basic research and is currently far removed from direct societal deployment. However, because our goal is to improve human–AI collaboration, caution is warranted when developing systems. Any future applications involving real users must prioritise safety, transparency, and accountability.
Acknowledgments
@@
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting C. Ruhdorfer. We especially thank R. Yang for implementing a prototype of the environment as part of her master’s thesis, and A. Penzkofer and L. Shi for numerous insightful discussions. J. Foerster is partially funded and J. Forkel is funded by the UKRI grant EP/Y028481/1 (originally selected for funding by the ERC).
Supplementary Materials
The following content was not necessarily subject to peer review.
Appendix A Servers and Software
@@
We ran our experiments on server systems running Ubuntu 24.04 equipped with NVIDIA Tesla V100-SXM2 GPUs with 32GB of memory, NVIDIA H100-NVL GPUs with 94GB of memory and a third server system featuring NVIDIA L40S with 48 GB of memory. All training runs are executed on a single GPU only. We trained our models using JAX (Bradbury et al., 2018), Flax (Heek et al., 2023) and Optax (DeepMind et al., 2020). In terms of tools, we did our analysis using NumPy (Harris et al., 2020), Pandas (Team, 2020), SciPy (Virtanen et al., 2020) and Matplotlib (Hunter, 2007). Our TransformerXL implementation is based on the one provided by Hamon (2024). Our single-file implementation of MAPPO is based on the one provided by JaxMARL (Rutherford et al., 2023) with additional inspiration taken from Hamon (2024).
Appendix B Accessibility and reusability
@@
The Yōkai Learning Environment (YLE) integrates directly into JaxMARL and implements its standard environment interface. As a result, any baseline method implementation of JaxMARL can be adapted to our environment quickly and with minimal effort. Upon acceptance, we will release the full environment implementation, training code for all reported baselines (including OP, HE-IPPO, and OBL), configuration files for reproducing the reported experiments, probing scripts, and pretrained model checkpoints. We have coordinated with the JaxMARL authors Rutherford et al. (2023), and upon acceptance YLE will be integrated into the official JaxMARL repository. This makes our open-source environment easily accessible to future researchers who will benefit from our reference implementations and those provided by the wider community. It also enables inter-environment comparison, facilitating broader benchmarking across cooperative multi-agent tasks. The project is accessible under https://git.hcics.simtech.uni-stuttgart.de/public-projects/Yokai_env.
Appendix C Fast, Scalable, and Parallel Yōkai in Graph-based JAX
@@ C.1 JAX-based Implementation
C.2 Reinforcement Learning in the YLE
@@
While in the original game, cards can be placed in any legal configuration, our environment limits the configuration to be in a predefined field of size $g\times g$ where $g=9$ in the 9C and $g=10$ in the 16C version. This enables the efficient computation of legal moves but prohibits some rare card configurations. While computing the exact number of disallowed configurations is infeasible, we find them to be extremely uncommon in practice.
For example, let’s assume a 2-player 9C game on a 9x9 grid. In such a setting, there a 4 hints and thus a total of 8 move actions for both players together. Cards are initially placed in a square in the centre of the board, leaving 3 empty spaces in the grid towards either side. In theory, thus, if agents agree to start placing cards in a row towards one of these sides, they would run out of space towards the border after 3 moves. In practice, this is seldom a suitable strategy and thus does not usually occur in games of Yōkai. This is because players want to group cards and are limited by the number of moves. They thus do not tend to start grouping cards in random directions, but rather group cards where they initially find a certain colour.
To verify that this is also true for our agents, we count the number of times when, in an evaluation game, an agent places a card on a border position. Note this is an overestimation as it does not account for the cases where agents place a card on the border but do not intend to place further cards in the same direction. We find that policies rarely to never use to border region of the board.
@@
For example, the action space is of size $1,068$ for 2-player 9C games and $2,322$ for the 16C version.
@@
Sclar et al. (2022) defines symmetric ToM environments as environments that (1) have a symmetric action space, (2) feature imperfect information, (3) allow the observation of others, and (4) require information-seeking behaviour. YLE features all four and thus qualifies. First, the action space is symmetric (see above). Second, information is imperfect since agents can never observe all cards first-hand and, for most of the game, only know a very limited subset of all cards. Third, it allows the observation of others both using explicitly encoded actions, as in ToMTrack or by observing their moves indirectly via their observations. Fourth, information needs to be acquired by observing cards first-hand or, second-hand, by interpreting moves and reasoning about knowledge. Agents that do not seek information and, for instance, repentantly observe the same card, will fail the game.
@@
Table 7: Number of hint cards according to the number of players and the environment version. The 9C version features three colours and thus has no three-colour hints.
16C 9C
@@
How many and which hint cards are available depend on the environment version and the number of players. For a 2-player game, there are 2 hints of one, 3 hints of two and 2 hints of three colours, i.e. 7 in total. Since there are 4 possible one-, 6 possible two- and 4 possible three-colour hints, players get a random subset of these hints at the start of the game per category. We give a full overview of all possible variations in Table 7. The number of hints directly influences the maximum number of turns in the game. This is because the game ends when all hints have been used. This means that games take longer if more players are added. Note, though, that this usually does not make the game easier. Instead, the number of cards one player can directly observe themself decreases with the number of players. For 16C YLE in the 2-player version, any agent can observe up to 14 of the 16 cars themselves (8 of 9 in 9C Yōkai). This shrinks to 10 of 16 for the 16C version and 6 of 9 for the 9C version when 4 players take part. Note how in the 3- and 4-player versions, agents are thus required to both work with less first-hand information and need to rely more on hints and interpreting the moves made by those around them. Not only this, but agents are tasked to do this tracking for more players as well. Thus, adding players to the YLE increases the difficulty substantially in two ways.
@@
C.3 Second Order ToM Reasoning Example Extended
Figure 8: Here we reexplain the second-order ToM reasoning example from the main paper over multiple timesteps in 2-player YLE in additional detail (see text): Refer to caption beliefs that Refer to caption beliefs that he (Refer to caption ) knew that cards 1 and 2 are of the same colour. Even though Refer to caption never saw card 1 and Refer to caption never observed card 3, both now know where all blue cards are, that they are grouped and that both share this knowledge as part of their common ground. In the future, they can now both observe and move different cards to finish early as they have one less card that needs to be observed.
Figure 8 displays the same example for second-order ToM reasoning in 2-player Yokai as the main paper. Here, we expand on the example. Figure 8 shows an agent ([Uncaptioned image] ) making inferences about what they assume the other agents ([Uncaptioned image] ) knowledge is due to their own previous action. In the example [Uncaptioned image] infers the colour of the card due to the behaviour of the other agent. [Uncaptioned image] reasons that because of how [Uncaptioned image] played, [Uncaptioned image] must have interpreted his move correctly. [Uncaptioned image] thus reasons that [Uncaptioned image] thinks that [Uncaptioned image] knew that cards 1 and 2 have the same colour and must go together. This is the case, [Uncaptioned image] made this inference correctly, and card 3 indeed is also blue. This reasoning allows agents to require less first-hand observations to get an overview of the board, potentially enabling them to end early.
@@ C.3 Second Order ToM Reasoning Example Extended
Appendix D Training Details
We share all the details of the training here, including all hyperparameters used in training and all details on the network architectures and how we tuned them. In general, we have tuned many aspects of our policy to arrive at the strongest agents we could for our experiments, see below. We took extra care in making these equal to make our results comparable.
D.1 Reinforcement Learning Details
Table 8: Hyperparameters of the learning process.
@@ D.1 Reinforcement Learning Details
Optimiser Adam (Kingma & Ba, 2015)
@@
Our work employs a standard IPPO algorithm (de Witt et al., 2020) for all experiments. We present all parameters of the learning process in Table 8. Only for OBL we reduce the number of parallel environments to $256$ since with $K=8$ we are otherwise unable to fit the OBL policies into memory. Note that this actually increases the number of gradient updates OBL policies recieve.
D.2 Neural Network Architectures
Our work employs several encoder architectures. We present all the details on them below.
General Policy Architecture
In general, in our work, no matter the encoder, the final policy network always has the same structure. That is given an embedding $e_{\text{encoder}}$ produced by an encoder network $f_{\text{encoder}}$, we pass the embedding to a linear projection, then a GRU (Cho et al., 2014) and finally either a to value or action head. We use ReLU (Fukushima, 1975; Nair & Hinton, 2010) as the activation function. The layers are initialised using orthogonal initialisation with a scale of $\sqrt{2}$ and bias 0. We have determined the number of policy layers and their size using hyperparameter tuning. Additionally, for all neural network architectures detailed below, we also varied and explored: The number of encoder blocks (2, 3 and 4), the hidden dimension of the encoder blocks, the aggregation function after the encoder (mean, sum, multiply and concatenate) and whether to anneal the dense rewards (False and True). We also compared addition additional hidden layers between GRU and the action/value heads (0, +1HL, +2HL). Finally, the value head produces a singular scalar output and is initialised using orthogonal initialisation with a scale of $0.00$ instead of $\sqrt{2}$. We initialise the action head in the same way but it produces logits for all possible actions instead.
GCN, GATv2 and the pre-embedding of Nodes